Express the number in normalized scientific notation:
0.003 x 10-2
Answer: 3.0 x 10^-5

Answer by Theo(13342)

your expression is:

.003 * 10^-2

the standard form would be 3.0 * 10 raised to an exponent.

in order to make 3.0 equivalent to .003, you have to divide it by 1000.

you get:

3 / 1000 = .003

this is the same as 3 * 1/1000 = .003

since 1/1000 equals 1/10^3 which is the same as 1 * 10^-3 which is the same as 10^-3, you get:

3 * 10^-3 = .003

your original expression is .003 * 10^-2

replace .003 with 3 * 10^-3 to get:

3 * 10^-3 * 10^-2 which is equal to 3 * 10^-3-2) which is equal to:

3 * 10^-5

you are correct.

confirm by using your calculator to get:

.003 * 10^-2 = .00003

3 * 10^-5 = .00003

actually my calculator said that .003 * 10^-2 = 3 * 10^-5.
it gave me the answer in scientific notation.
calculator do that sometimes.